Circular No. 7142

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1999bt IN ANONYMOUS GALAXY G. M. Hurst, Basingstoke, England, reports the discovery by T. Boles,

Wellingborough, of an apparent supernova (mag ∼ 17) on exposures taken with a 0.14-m f/11 (+ AP7 camera) for the U.K. Nova/Supernova Patrol on Apr. 9.120 and 13.088 UT. The new object is in the field of IC 1254 but is probably linked to a much fainter galaxy just south of the suspect. M. Armstrong, Rolvenden, has measured the following precise position from the discovery exposure of Apr. 9: α = 17^{h}11^{m}37^{s}.77, δ = +72^{o}25′15′′.5 (equinox 2000.0). Hurst notes that the new object does not appear on the Palomar Sky Survey (1953 June 15) or on the second generation sky survey images (epoch 1993.625) at the Space Telescope Science Institute website.

S. Jha, P. Garnavich, P. Challis, and R. Kirshner, Harvard-Smithson-ian Center for Astrophysics, report that a spectrum of SN 1999bt was obtained by M. Calkins on Apr. 14.5 UT with the F. L. Whipple Observatory (FLWO) 1.5-m telescope (+ FAST spectrograph) and reveals it to be a type-Ia supernova, ∼ two weeks past maximum light. Concurrent images taken at the FLWO 1.2-m telescope (+ 4Shooter) by L. Macri yield V = 18.1 for the supernova, which is located 4′′ west and 10′′ north of the host galaxy. Cross-correlation of the host-galaxy spectrum with an absorptionline template yields a recession velocity of 14 960 ± 60 km/s.

COMET C/1999 G2 (SOHO) D. A. Biesecker, NASA Goddard Space Flight Center, reports obser-

vations of a probable Kreutz sungrazing comet in SOHO/LASCO C2 data on Apr. 13. The comet reached a peak apparent magnitude of ∼ 6.5, and a very short, faint tail is evident. It was not visible in the C3 coronagraph because of the vignetting caused by the pylon that holds the solar occulter.

1999 UT  α_{2000}  δ_{2000}
Apr. 13.338  1^{h}32.^{m}0  + 7^{o}58′

Astrometric measurements, made by Biesecker and reduced by B. G. Marsden, and orbital elements by Marsden, are given on MPEC 1999-G25.

SUPERNOVA 1999ac IN NGC 6063 Unfiltered CCD magnitude estimates: Feb. 28.111 UT, 15.3 ± 0.3 (L.

Kiss and K. Sárneczky, Szeged, Hungary); Mar. 3.726, 15.7 (S. Yoshida and K. Kadota, Ageo City, Japan); 12.803, 14.8 (Yoshida); 21.13, 14.7 (S. Moretti, Forlí, Italy); 31.753, 15.5 (Yoshida); Apr. 4.751, 15.5 (Yoshida).

1999 April 14  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green